# WCCNet: Wavelet-context Cooperative Network for Efficient Multispectral Pedestrian Detection

### v2

@@ Abstract
Multispectral pedestrian detection achieves better visibility in challenging conditions and thus is essential to autonomous driving, for which both the accuracy and computational cost are of paramount importance. Most existing approaches treat RGB and infrared modalities equally. They typically adopt two symmetrical backbones for multimodal feature extraction, which ignore the substantial differences between modalities and bring great difficulty for the reduction of the computational cost as well as effective crossmodal fusion. In this work, we propose a novel and efficient framework named Wavelet-context Cooperative Network (WCCNet) that is able to differentially extract complementary features of different spectra with lower computational complexity, and further fuse these diverse features based on their spatially relevant crossmodal semantics. In particular, WCCNet simultaneously explore wavelet context and RGB textures within a cooperative dual-stream backbone, which is composed of adaptive discrete wavelet transform (ADWT) layers and heavyweight neural layers. The ADWT layers extract frequency components for infrared modality, while neural layers handle RGB modality features. Since ADWT layers are lightweight and extract complementary features, this cooperative structure not only significantly reduces the computational complexity, but also facilitates the subsequent crossmodal fusion. To further fuse these infrared and RGB features with significant semantic differences, we elaborately design the crossmodal rearranging fusion module (CMRF), which can mitigate spatial misalignment and merge semantically complementary features in spatially-related local regions to amplify the crossmodal reciprocal information. Results from comprehensive evaluations on KAIST and FLIR benchmarks indicate that WCCNet outperforms state-of-the-art methods with considerable computational efficiency and competitive accuracy simultaneously.
I Introduction
Pedestrian detection[1, 2, 3] is a crucial task widely studied in autonomous driving field owing to its central role in ensuring safety and preventing accidents[4]. Due to the limitations of RGB cameras in capturing pedestrian features under low-light environments, the integration of infrared modality with RGB modality considerably enhance the reliability[5, 6, 7], and thus has attracted growing research interests. Recent years have witnessed rapid development of deep learning-based multispectral pedestrian detection methods.
@@ I Introduction
Figure 1: Inference time (ms) vs. $\mathrm{MR}^{-2}$ comparisons with state-of-the-art methods on the KAIST[5] test set, including MSDS[8], ARCNN[9], CIAN[10], TC-DET[11], MBNet[12], MLPD[13], BAANet[14], CMPD[15], and TINet[16]. WCCNet-L, WCCNet-S, and WCCNet-XS are variants of WCCNet with different amounts of parameters. A lower $\mathrm{MR}^{-2}$ represents higher precision. In this way, methods located at lower left corner possess better performance.
@@
Existing multispectral detection methods[17, 9, 18, 8, 13, 12, 19, 20, 16, 11, 14, 15, 10] equally treat the feature extraction of RGB and infrared modalities, and typically apply symmetrical backbones for both RGB and infrared modalities, which will inevitably entail extra computational burden. Therefore, it is quite significant and challenging to find a new network with fewer parameters and high accuracy simultaneously. On the other hand, infrared images and RGB images have clearly distinct features[14, 15]. While RGB images have richer information about textures and colors, infrared images contain valuable geometry details such as shapes, contours, and areas, which will be manifested in high-frequency components. Thus, we explore the integration of wavelets and neural layers to differentially extract features from different modalities. This is inspired by the fact that wavelet-based networks can achieve superior performance in various image applications involving extraction of different frequency features [21, 22, 23].
@@
Furthermore, the obvious visual differences between RGB and infrared modalities also signify a huge semantic distinction between their extracted features, which has posed a challenge to crossmodal fusion. Conventional fusion schemes[18, 8, 9, 13] directly apply simple pixel-wise addition and channel-dimension concatenation, thus they are weak in extracting complementary features. Recent fusion schemes [12, 19, 20, 16] differentiate complementary features from redundant features along channel dimension based on the differences in their numeric values. They assume explicitly or implicitly that features of different modalities are numerically homogeneous. However, it is well-known that the semantic differences between features of different modalities cannot be adequately represented by their numeric differences, especially when there exists a huge semantic gap between them. Since pedestrians in the paired RGB and infrared images are spatially related, we manage to explore semantic correlation between crossmodal features based on their spatial correlation.
@@
In this paper, aiming for faster and more accurate multispectral pedestrian detection, we propose a novel Wavelet-Context Cooperative framework (WCCNet), consists of ADWT-integrated dual-stream backbone and crossmodal rearranging fusion module. The proposed WCCNet differentially extracts features for infrared and RGB modalities with lower computational cost, and effectively aggregates semantically complementary features through the crossmodal context in spatially relevant local regions.
Firstly, we propose a dual-stream backbone composing of CNN layers and adaptive DWT (ADWT) layers, where CNN layers extract spatial-domain features from RGB modality and ADWT layers extract frequency features from infrared modality. The reciprocal relationship between CNN and ADWT layers not only significantly reduces the amount of parameters and inference time, but also improves the extraction of infrared features to complement RGB modality and facilitate the subsequent crossmodal fusion.
@@
Our main contributions are summarized as follows:
  • A novel framework WCCNet is proposed to achieve considerable detection efficiency and competitive accuracy. The ADWT and neural layers of its dual-stream backbone possess a reciprocal relationship, which conduct differential manners for different modalities to extract complementary features and prepare for further fusion.
  • CMRF is proposed to fuse semantically complementary and spatially relevant features of the two modalities with large discrepancy. Cross-domain Embedding (CE) and Crossmodal Spatial Alignment (CSA) layers are integrated into CMRF module to tackle the cross-domain semantic gap and the crossmodal spatial misalignment problem respectively.
@@
II Related Work
Multispectral Pedestrian Detection
@@ II Related Work
Multispectral pedestrian detection has attracted a growing research interest for its essential role in autonomous driving. In the past few years, deep learning-based methods[17, 9, 18, 8, 13, 12, 19, 20, 16, 11, 14, 15, 10] have exhibited state-of-the-art performance and become the mainstream. Most of them typically utilize a dual-stream framework to extract RGB and infrared features both by symmetrical neural layers, while focussing on the multimodal feature extraction and crossmodal feature fusion. Liu et al. [17] employed four fusion architectures which fuse at different stages of feature extraction, and found that the middle fusion produced the highest accuracy. Based on this observation, MSDS[8] further improved the detection accuracy by refining the detection results using cascaded classification subnetwork. Considering position shift problem in weakly aligned multispectral images, AR-CNN[9] designed an aligned region network to adaptively align the multimodal features. Later, for totally unpaired multispectral images, MLPD[13] leveraged multi-label learning to obtain competitive detection accuracy in unpaired conditions. As for crossmodal feature fusion, beyond conventional fusion methods including pixel-wise addition[9] and channel-dimension concatenation[18, 8, 13], MBNet[12] presented a differential modality aware fusion (DMAF) module for better complementary feature extraction. Following the DMAF paradigm of MBNet, there witnessed a series of similar methods that improved upon MBNet and achieved state-of-the-art performance, e.g., PIAFusion[19], RISNet[20], and TINet[16], etc. In addition to feature-level fusion, CMPD[15] conducted decision-level fusion based on Dempster’s combination rule. YOLO-MSLite[7] further explores to improve detection efficiency.
@@
However, these prior works have neglected the intrinsic feature differences between RGB and infrared modalities, and simply utilized symmetrical backbones for RGB and infrared feature extraction. In fact, for infrared modality rich in contour features but deficient in texture features, it is unnecessary to apply heavyweight neural layers at shallow stages, which may results in excessive computational burden and even feature redundancy. Instead, we explore to extract features differentially according to the characteristics of different modalities with lower computational burden.
@@
Meanwhile, crossmodal fusion schemes in prior works [12, 19, 20, 16] assumed features of different spectra are numerically homogeneous. Accordingly, they tended to filter out redundant features with minor numeric differences from multispectral features. However, due to the huge semantic gap between modalities, redundant features of different modalities can present a large discrepancy in their numeric values, which will hinder the crossmodal fusion of these methods. Therefore, in this study, we establish the semantic-aware interaction between crossmodal features based on their spatial consistency, and further exploit their complementary information.
Wavelets in Deep Learning
@@
As an effective tool in time-frequency analysis, wavelets have been widely applied in various signal processing applications [24, 25]. A great deal of researches has been devoted to the combination of fixed wavelets and deep learning networks for both low-level and high-level visual tasks. All these works demonstrate that superior results can be achieved by using wavelet-based deep learning methods. It is natural yet quite non-trivial to design networks learning directly from frequency domain via wavelet transforms on input data [23]. In wavelet scattering networks, the wavelet transform convolutions are cascaded with nonlinear activations and averaging operators [22]. Fujieda et al. [26] proposed a wavelet CNN structure, in which the Haar wavelets introduce multiscale frequency components into network at different stages as a supplement to multiresolution analysis. To suppress the aliasing effect, Li et al. [27] proposed a novel CNN framework named WaveCNet by replacing of the down-sampling operations with fixed DWTs, achieving higher classification accuracy and better noise-robustness. For semantic segmentation of biomedical images, Zhou et al. [28] utilized wavelets to alleviate the learning bias by decomposing images into low-frequency and high-frequency parts.
Recently, there are some researches about fully learnable wavelets[21, 29]. Differing from these literatures, with the goal to prioritize computational efficiency, ADWT layers of WCCNet adopt fixed Haar wavelets modulated by only a few learnable gain coefficients. While fixed Haar wavelets provide computational simplicity, the learnable gain coefficients can strengthen useful frequency components. Thereby WCCNet achieves an optimal balance between accuracy and efficiency. Besides, instead of stacking wavelet with CNN layers in a cascade manner, in which DWT proceeds or follows other types of learning networks [23, 27, 30], WCCNet combines different layers in a parallel and cooperative way. Reciprocal relationship between CNN-extracted RGB features and ADWT-extracted infrared features is further exploited to improve detection accuracy and speed.
Figure 2: Overview framework of WCCNet. The WCCNet consists of three parts: mutli-domain feature extraction backbone, crossmodal rearranging fusion (CMRF) module and detection head. In backbone, WCCNet utilize CNN layers and ADWT layers to extract features of RGB and infrared modalities respectively. CMRF fusion modules are embedded into the first three stages to supplement crossmodal complementary information. The multiscale features of the last three stages are aggregated by feature pyramid mechanism[31], and then are decoded by decoupling detection head[32] to generate final predictions.
Multispectral Fusion Scheme
@@
Fusion between feature maps of different semantic contents is essential for multispectral detection task. Conventional fusion schemes regard input features as a whole and conduct simple numeric combinations, e.g., pixel-wise addition, pixel-wise multiplication, and channel-dimension concatenation, which are incapable of highlighting the crucial parts of input features. In recent years, fusion schemes based on attention mechanisms have emerged. SENet[33] applied channel attention mechanism and boosted its accuracy by weighting each channel based on channel-wise spatial pooling results. Based on SENet, CBAM[34] additionally adopted spatial attention mechanism to spatially weight feature maps with a score map calculated from pixel-wise channel pooling. The variants of the aforementioned channel attention and spatial attention mechanisms were widely used as the fusion schemes of prior multispectral detection works[12, 11, 14, 19, 20, 16].
@@
Recently, there are some works about learning dynamic kernels which may be useful for fusion tasks. CARAFE++[35] and IndexNet[36] were proposed to generate adaptive kernels to aggregate contextual information within related receptive filed. However, for multispectral features with a large semantic gap, these single-modal fusion schemes need pre-concatenation for multimodal input features, and thus can not capture contextual correlations between feature maps of different modalities. Besides, these fusion schemes heavily rely on pixel-wise or channel-wise alignment of input features, thus prone to feature misalignment which are common in practical applications. Instead, the contextual correlations between multimodal features of different modalities and the adaptive spatial feature alignment are explored in this work.
III Proposed Method

### v3

@@ Abstract
Multispectral pedestrian detection is essential to various tasks especially autonomous driving, for which both the accuracy and computational cost are of paramount importance. Most existing approaches treat RGB and infrared modalities equally. They typically adopt two symmetrical backbones for multimodal feature extraction, which ignore the substantial differences between modalities and bring great difficulty for the reduction of the computational cost as well as effective crossmodal fusion. In this work, we propose a novel and efficient framework named Wavelet-context Cooperative Network (WCCNet), which differentially extracts complementary features across spectra with low computational cost and further fuses these diverse features based on their spatially relevant cross-modal semantics. WCCNet explores an asymmetric but cooperative dual-stream backbone, in which WCCNet utilizes generic neural layers for texture-rich feature extraction from RGB modality, while proposing Mixture of Wavelet Experts (MoWE) to capture complementary frequency patterns of infrared modality. By assessing multispectral environmental context, MoWE generates routing scores to selectively activate specific learnable Adaptive DWT (ADWT) layers, alongside shared static DWT, which are both considerible lightwight and efficient. This cooperative structure not only significantly reduces the computational complexity, but also facilitates the subsequent crossmodal fusion. To further fuse these multispectral features with significant semantic differences, we elaborately design the crossmodal rearranging fusion module (CMRF), which aims to mitigate misalignment and merge semantically complementary features in spatially-related local regions to amplify the crossmodal reciprocal information. Results from comprehensive evaluations on KAIST and FLIR benchmarks indicate that WCCNet outperforms state-of-the-art methods with considerable computational efficiency and competitive accuracy.
keywords:
Multispectral Pedestrian Detection , Wavelet-context Cooperation , Mixture of Wavelet Experts , Crossmodal Rearranging Fusion
journal: arXiv
[…]
1 Introduction
Pedestrian detection is a crucial task widely studied for autonomous driving owing to its central role in ensuring safety and preventing accidents. Due to the limitations of RGB cameras in capturing pedestrian features under low-light environments, the integration of infrared modality with RGB modality [MPD-1, MPD-2] considerably enhances the reliability [KAIST2015, FLIRDataset], and attracts growing research interests. Recent years have witnessed rapid development of deep learning-based multispectral pedestrian detection methods.
@@ 1 Introduction
Existing multispectral detection methods [MSDS2018, ARCNN2019, CIAN2019, MBNet2020, MLPD2021, CMPD2022, liu2016CleanKAIST, RPN2017Konig, tang2022piafusion, wang2022RISNet, zhang2023TripleNet, TCDET2020, yang2022BAANet, ICAFusion] equally treat the feature extraction of RGB and infrared modalities, and typically apply symmetrical backbones for both RGB and infrared modalities. This symmetry ignores the inherent disparity between spectra, and will inevitably entail extra computational burden. Therefore, it is quite significant and challenging to find a new network with fewer parameters and high accuracy simultaneously. On the other hand, infrared images and RGB images have clearly distinct features [CMPD2022, yang2022BAANet]. RGB images have richer information about textures and colors [MD-1], which necessitate dense network to resolve. Conversely, infrared images lack internal texture, and mainly contain valuable geometry details including low-frequency thermal blobs and high-frequency contours. Processing infrared stream with heavyweight neural layers results in significant computational redundancy, and may overfit to thermal sensor noise like quantization artifacts. It is suggested to design dynamic frequency-domain filters for efficient infrared feature extraction.
Figure 1: Inference time (ms) vs. $\mathrm{MR}^{-2}$ comparisons with state-of-the-art methods on the KAIST [KAIST2015] test set. MSDS [MSDS2018], ARCNN [ARCNN2019], CIAN [CIAN2019], MBNet [MBNet2020], MLPD [MLPD2021], CMPD [CMPD2022], GAFF [GAFF2021], and MS-DETR [MS-DETR2024] are involved. WCCNet, WCCNet-S, and WCCNet-XS are three variants with different amounts of parameters. Lower $\mathrm{MR}^{-2}$ represents higher precision, lower inference time denotes higher detection speed.
@@
Furthermore, the obvious visual differences between RGB and infrared modalities also signify a huge semantic distinction between their extracted features, which has posed a challenge to crossmodal fusion. Conventional fusion schemes [MSDS2018, ARCNN2019, MLPD2021, RPN2017Konig] directly apply simple pixel-wise addition and channel-dimension concatenation, thus they are weak in extracting complementary features. Recent fusion schemes [MBNet2020, tang2022piafusion, wang2022RISNet, zhang2023TripleNet] differentiate complementary features from redundant features along channel dimension based on the differences in their numeric values. They assume explicitly or implicitly that features of different modalities are numerically homogeneous. However, it is well-known that the semantic differences between features of different modalities cannot be adequately represented by their numeric differences, especially when there exists a huge semantic gap between them. Since pedestrians in the paired RGB and infrared images are spatially related, we manage to explore semantic correlation between crossmodal features based on their inherent correlation.
@@
To tackle with abovementioned limitations, we propose Wavelet-Context Cooperative framework (WCCNet) for faster and more accurate multispectral pedestrian detection, consisting of MoWE-integrated cooperative backbone and crossmodal rearranging fusion module. WCCNet asymmetrically extracts features for infrared and RGB modalities with lower computational cost, and effectively aggregates semantically complementary features through crossmodal context in spatially relevant local regions.
Figure 2: Feature visualization of WCCNet on KAIST testing set. For complex illumination scenarios, WCCNet effectively captures and integrates complementary features by wavelet-context cooperative framework.
Firstly, we propose a wavelet-cooperative backbone that combines Mixture of Wavelet Experts (MoWE)–based infrared branch with generic neural layer-based RGB branch. MoWE is designed to route channel-wise features to a diverse bank of wavelet filters, including a global context-aware router and two kinds of wavelet experts. By assessing environmental context jointly from multispectral inputs, MoWE first produces routing scores that selectively activate a sparse subset of routed experts. We propose Adaptive Discrete Wavelet Transform (ADWT) layers as routed experts, alongside with static DWT layers as shared experts that are always activated. ADWT experts are constructed as learnable parallel DWT layers with adaptive weighted scores, allowing optimization of wavelet basis to better capture pedestrian-specific frequency patterns. Complementing these are shared DWT experts utilizing pre-defined static weights as stable frequency prior. MoWE is lightweight compared to generic neural layers, and its sparse activation further enhances inference efficiency. As for RGB branch, we pratically employ convolutional neural network (CNN) as generic neural layers. WCCNet framework also supports transformer-based layers as generic neural layers, e.g., ViT and Swin Transformer, but have not been explored in this paper. The reciprocal relationship between MoWE and generic neural layers not only significantly reduces computational cost, but also yields informative infrared features to complement RGB modality and facilitate the subsequent crossmodal fusion, as shown in Fig. 2.
@@
The framework of WCCNet is illustrated in Fig. 3 and Fig. 4, and our main contributions are summarized as follows:
  • 1.
    A novel framework WCCNet is proposed to achieve considerable detection efficiency and competitive accuracy. WCCNet employs an asymmetric wavelet-cooperative backbone, utilizing lightweight MoWE to dynamically routes infrared inputs through activated ADWT experts to capture informative frequency-domain features, while extracting RGB features by generic neural layers. It leverages reciprocal relationship between MoWE and generic neural layers to extract complementary features with lower computational cost.
  • 2.
    CMRF is proposed to fuse semantically complementary and spatially relevant features of the two modalities with large discrepancy, generating on-the-fly attention weights to focus on spatially related and semantically complementary regions. Cross-domain Embedding (CE) and Crossmodal Spatial Alignment (CSA) layers are integrated into CMRF module to tackle the cross-domain semantic gap and the crossmodal spatial misalignment respectively.
  • 3.
@@
2 Related Work
2.1 Multispectral Pedestrian Detection
@@
Multispectral pedestrian detection has attracted a growing research interest for its essential role in autonomous driving. In the past few years, deep learning-based methods [MSDS2018, ARCNN2019, CIAN2019, MBNet2020, MLPD2021, CMPD2022, liu2016CleanKAIST, RPN2017Konig, tang2022piafusion, wang2022RISNet, zhang2023TripleNet, TCDET2020, yang2022BAANet] have exhibited state-of-the-art performance and become the mainstream. Most of them typically utilize a dual-stream framework to extract RGB and infrared features both by symmetrical neural layers. Liu et al. [liu2016CleanKAIST] employed four fusion architectures which fuse at different stages of feature extraction, and found that the middle fusion produced the highest accuracy. Based on this observation, MSDS [MSDS2018] further improved the detection accuracy by refining the detection results using cascaded classification subnetwork. Considering totally unpaired multispectral images, MLPD[MLPD2021] leveraged multi-label learning to obtain competitive detection accuracy in unpaired conditions. As for crossmodal feature fusion, beyond conventional fusion methods including pixel-wise addition [ARCNN2019] and channel-dimension concatenation [MSDS2018, MLPD2021, RPN2017Konig], MBNet [MBNet2020] presented a differential modality aware fusion (DMAF) module for better complementary feature extraction. Following the DMAF paradigm of MBNet, there witnessed a series of similar methods that achieved state-of-the-art performance, e.g., PIAFusion [tang2022piafusion], RISNet [wang2022RISNet], and TINet [zhang2023TripleNet], etc. Employing an iterative cross-attention mechanism, ICAFusion [ICAFusion] progressively refine and fuse complementary multispectral features. Besides, MS-DETR [MS-DETR2024] explored loosely coupled fusion strategy within a multi-modal transformer decoder. In addition to feature-level fusion, CMPD [CMPD2022] conducted decision-level fusion based on Dempster’s combination rule.
@@
However, these prior works have neglected the intrinsic feature differences between RGB and infrared modalities, and simply utilized symmetrical backbones for both modalities. In fact, for infrared modality rich in contour features but deficient in texture features, it is unnecessary to apply heavyweight neural layers at shallow stages, which may results in excessive computational burden and even feature redundancy. Instead, we explore to extract features differentially according to the characteristics of different modalities with lower computational burden.
@@
Meanwhile, crossmodal fusion schemes in prior works [MBNet2020, tang2022piafusion, wang2022RISNet, zhang2023TripleNet] assumed features of different spectra are numerically homogeneous. Accordingly, they tended to filter out redundant features with minor numeric differences from multispectral features. However, due to the huge semantic gap between modalities, redundant features of different modalities can present a large discrepancy in their numeric values, which will hinder the crossmodal fusion of these methods. Therefore, in this study, we establish the semantic-aware interaction between crossmodal features based on their spatial consistency, and further exploit their complementary information.
2.2 Wavelets in Deep Learning
@@
As an effective tool in time-frequency analysis, wavelets have been widely applied in various signal processing applications [mallat1989wavelet, mallatWaveletBook]. A great deal of research has been devoted to the combination of wavelet and deep learning networks for both low-level and high-level visual tasks, demonstrating that superior results can be achieved by such methods. Early works primarily explored the integration of fixed wavelet transforms into deep networks. It is natural to design networks learning directly from the frequency domain [deepinwavelet]. Further, in this context, wavelet scattering network utilizes fixed wavelet transform convolutions to cascade with nonlinear activations and averaging operators [MallatDWT]. Fujieda et al. [WaveletConvNet] proposed a structure to introduce multiscale frequency components at different stages, while Li et al. [li2021wavecnet] proposed WaveCNet to suppress aliasing effects by replacing down-sampling operations with fixed DWTs, achieving better noise-robustness. More recently, research attention has shifted towards fully learnable wavelets. Michau et al. [GabrielPNAS] construct learnable wavelet following conjugate quadrature filter bank property, showing effectiveness for unsupervised monitoring of high-frequency time series. Furthermore, the adaptive wavelet distillation (AWD) method [WaveletDistill] learns an effective wavelet transform by penalizing feature attributions in the wavelet domain, yielding SOTA performance for cosmology inferring and molecular-partner prediction. Motivated by inherent high efficiency of wavelet transform, we further explore a dynamic mixture mechanism of fixed and learnable wavelets to facilitate the extraction of diverse frequency characteristics for infrared modality while maintaining computational economy.
2.3 Multispectral Fusion Scheme
@@
Fusion between feature maps of different semantic contents is essential for multispectral detection task. Conventional fusion schemes regard input features as a whole and conduct simple numeric combinations, e.g., pixel-wise addition, pixel-wise multiplication, and channel-dimension concatenation, which are incapable of highlighting the crucial parts of input features. In recent years, fusion schemes based on attention mechanisms have emerged. SENet [hu2018squeeze] applied channel attention with channel-wise weights from spatial global information, while CBAM [woo2018cbam] additionally adopted spatial attention with a score map calculated from pixel-wise channel pooling. The variants of the aforementioned channel and spatial attention mechanisms were widely used as the fusion schemes of prior multispectral detection works [MBNet2020, tang2022piafusion, wang2022RISNet, zhang2023TripleNet, TCDET2020, yang2022BAANet].
@@
Recently, there are some works about learning dynamic kernels which may be useful for fusion tasks. CARAFE++ [wang2021carafe++] and IndexNet [lu2022index] were proposed to generate adaptive kernels to aggregate contextual information within related receptive filed. However, for multispectral features with a large semantic gap, these single-modal fusion schemes need pre-concatenation for multimodal input features, and thus can not capture contextual correlations between feature maps of different modalities. Besides, these fusion schemes heavily rely on pixel-wise or channel-wise alignment of input features, thus prone to feature misalignment which are common in practical applications. Instead, the contextual correlations between multimodal features of different modalities and the adaptive spatial feature alignment are explored in this work.
3 Proposed Method
3.1 MoWE-integrated Wavelet-cooperative Backbone
3.1.1 Overall Structure
For higher computational efficiency and precision, we propose an asymmetric wavelet-cooperative backbone that extracts features differentially according to the characteristics of different modalities, as shown in Fig. 3. It combines a Mixture of Wavelet Experts (MoWE)–based infrared branch with a generic neural layer-based RGB branch. MoWE has fewer parameters and higher computational efficiency than generic neural layers. Compared with conventional symmetric designs, the integration of MoWE substantially enhances the efficiency of asymmetric backbone.